OpenStudy (anonymous):

(2xy+3)dx+(x^2-1)dy=0 exact equations I get that they are exact and get to f(x,y)=Integral(2xy+3)dx not sure if it should equal x^2+3x+g(y) or x^2+g(y)

OpenStudy (anonymous):

im gonna go with \[x^2+3x+g(y)\]

OpenStudy (anonymous):

So you're looking for an intermediate step? Or do you think that these are your final solutions?

OpenStudy (anonymous):

no i know that these aren't our final solutions. you continue to \[x^2+3x+g'(y)=X^2-1\] \[g'(y)=-3x-1\] \[g(y)=(-3/2) x^2-1x\] then i stopped cause i don't think this is right:/

OpenStudy (anonymous):

\[\Large f_x+f_y\frac{dy}{dx}=0 \] where f(x,y) is multivariable. You checked the exactness for this equation already. so you want to solve either: \[\Large f_x=2xy+3 \] or \[\Large f_y=x^2-1 \] Lets choose the first one: \[ \Large f_x=\frac{\partial f(x,y)}{\partial x}=2xy+3\] and integrate with respect to x to get one representation for f(x,y) \[\Large f(x,y)=\int\partial f(x,y)=x^2y+3x+g(y)+k \]

OpenStudy (anonymous):

ok so i missed the y in the x^2y. so now we \[x^2y+3x+g'(y)=x^2-1\] ??

OpenStudy (anonymous):

differentiate your temporary expression f(x,y) with respect to partial y. \[\Large f_y=\frac{\partial f(x,y)}{\partial y}=x^2+g'(y)=x^2-1 \] Simplify an integrate: \[\Large g'(y)=-1 \\ \Large \therefore g(y)=-y\]

OpenStudy (anonymous):

Awesome! thank you so much!

OpenStudy (anonymous):

you're very welcome :-)